# A Unified MDL-based Binning and Tensor Factorization Framework for PDF Estimation

### v1

@@ IV-A Synthetic Data
First, we investigate the impact of candidate cut strategies on the performance of the MDL algorithm. We consider a univariate Gaussian mixture with $R=6$ components. We estimate the histogram using $T=20,\!000$ samples and the results are averaged over 50 independent trials. Table I summarizes the performance of our proposed quantile-based approach compared with the two-cuts and midpoint approaches in [8]. We evaluate the performance in terms of the average number of bins, the MDL score, the runtime (in minutes), and the negative log-likelihood (NLL). The quantile-based method estimates an average of 20.08 bins with the lowest MDL score of $59,\!397$ and an average runtime of only 1.47 minutes, while achieving an NLL of $13,\!139$ (lower is better). In contrast, the two-cuts approach produces slightly fewer bins (19.02) and a comparable NLL ($13,\!139$) but at a substantially higher runtime of 71.32 minutes. The midpoints approach, with an average of 19.3 bins and an NLL of $13,\!141$, requires 15.83 minutes. These results indicate that restricting the candidate set via quantiles does not impose a significant loss in optimality, but drastically reduces runtime.
Next, we evaluate our MDL-based PDF estimation algorithm against a uniform discretization baseline using a simulation setup similar to [6]. Here, data are generated from an $N=5$-dimensional Gaussian mixture of $R=6$ components, and the accuracy of the model is measured by the KLD between the true and estimated densities computed by Monte Carlo integration averaged over 100 independent trials. We compare the performance of our algorithm to that of the classical Gaussian mixture model based on EM (EM GMM), coupled tensor factorization algorithm based on alternating optimization and a KLD loss criterion (CTF-AO-KL) [6], and the “Oracle” method, which assumes that the labels (latent states) are known and serves as an empirical lower bound for the KLD. As shown in Fig. 3, our approach outperforms the uniformly binned (20 bins) CTF-AO-KL in both KLD and clustering accuracy. This result underscores the effectiveness of MDL-based binning in adapting to skewed and non-uniform densities, whereas uniform binning–due to its fixed allocation– often leads to a model mismatch when the data deviate from uniformity. Although EM GMM remains a practical choice for modeling Gaussian mixtures, its performance can be compromised if the discretization is not well adapted to the underlying density.
@@ IV-B Real Data
TABLE I: Comparison Between Different Cut-Point Placement Strategies
Method Avg. No. of Bins MDL Score Runtime [min] NLL
Proposed 20.08 59397 1.4735 13139.14
Two-cuts 19.02 59423 71.3157 13138.88
Midpoints 19.30 59417 15.8292 13141.42
We further evaluate our proposed approach on a real dry bean dataset [2]. In this experiment, each sample is characterized by features from 7 different dry bean varieties, and the objective is to accurately predict the bean class. Initially, we estimate the model rank using variational Bayesian inference (VB-PMF) [20]; given an upper bound for the rank $R$, this algorithm automatically prunes irrelevant components to determine the effective rank. The estimated rank ($R=48$ the most common) is then used to train our method and other competing algorithms. We randomly split the data into training set ($80\,\%$) and testing set ($20\,\%$) and evaluate the model’s performance using the testing set. We calculate the classification accuracy (defined as the proportion of correctly classified samples to the total number of samples) and report the results averaged over 50 random data splits.
In general, our method achieves higher classification accuracy while also significantly reducing computational time compared to the baselines.
@@ V Conclusion
In this paper, we introduced a unified framework that combines MDL and tensor factorization for non-parametric PDF estimation. By integrating a quantile-based cutting strategy, we improved the MDL computation time without compromising the performance. Experimental results on both synthetic and real datasets demonstrated significant improvements in estimation accuracy and computational efficiency compared to conventional uniform binning methods.
TABLE II: Comparison of Models Based on Classification Accuracy (Class. Acc.) and Runtime

### v2

@@ IV-A Synthetic Data
We first investigate how the candidate-cutting strategy affects the MDL algorithm. A six-component univariate Gaussian mixture is sampled with T=20  [^0] T=20\,000 observations, the experiment is repeated 50 times. Each method starts with the same upper limit of $K_{\text{max}}=50$ bins and searches for the MDL-optimal bins. Fig. 4 displays box-plots for the proposed quantile method, the two-cuts, and mid-points heuristics of [7]. We report (i) the bins count, (ii) the MDL score  [^3] We refer to (6) as the MDL score which quantifies the quality of the MDL histogram., (iii) the runtime (in minutes), and (iv) the negative log-likelihood (NLL). The MDL-score and NLL box-plots almost entirely overlap across the three strategies; medians and inter-quartile ranges coincide, indicating statistically similar fit quality. The runtime boxes, however, are spread over two orders of magnitude: the quantile method finishes in under 1 min, the mid-points grid centers around 20 min, and the two-cuts exceeds 70 min. These results indicate that restricting the candidate set via quantiles does not impose a significant loss in MDL optimality or likelihood fit, yet drastically improve computational efficiency.
Next, we evaluate our MDL-based PDF framework against a uniform discretization baseline following the set-up of [5]. Here, data are generated from five-dimensional Gaussian mixture ($N=5$, $R=6$), and the accuracy is measured by the KLD between the true and estimated densities, averaged over 100 Monte-Carlo trials. We compare the performance of our algorithm to that of the classical Gaussian mixture model based on EM (EM GMM), coupled tensor factorization algorithm based on alternating optimization and a KLD loss criterion (CTF-AO-KL) [5], and the “Oracle” method, which assumes that the labels (latent states) are known and serves as an empirical lower bound for the KLD. As shown in Fig. 3 (a),(b), our approach outperforms the uniformly binned (20 bins) CTF-AO-KL in both KLD and clustering accuracy. Fig. 3 (c) demonstrates how the histogram grows with $T$. When $T$ is small, adding many cuts would leave only a few points per bin, so the NML term cannot compensate for the MDL penalty. As more data become available, each prospective bin contains enough points to estimate its probability reliably; the NML term now rewards a finer partition, making higher-resolution histogram both feasible and favorable. Although the EM GMM remains a practical choice for modeling Gaussian mixtures, its performance can be compromised if the discretization is not well adapted to the underlying density.
@@ IV-B Real Data
Figure 4: Performance of Cuts Placement Strategies.
We further evaluate our proposed approach on a real dry bean dataset [19]. In this experiment, each sample is characterized by features from 7 different dry bean varieties, and the objective is to accurately predict the bean class. Initially, we estimate the model rank using variational Bayesian inference (VB-PMF) [20]; given an upper bound for the rank $R$ (e.g. the maximum rank for which the Kruskal identifiability condition for the CPD is satisfied [10]), this algorithm automatically prunes irrelevant components to determine the effective rank. The estimated rank ($R=48$) is then used to train our method and other competing algorithms. We randomly split the data into training set ($80\,\%$) and testing set ($20\,\%$). We calculate the classification accuracy (defined as the proportion of correctly classified samples to the total number of samples) and report the results averaged over 50 random data splits. In general, as demonstrated in Table I, our method achieves higher classification accuracy while also significantly reducing the computational cost.
@@ V Conclusion
In this paper, we introduced a unified framework that combines MDL and tensor factorization for non-parametric PDF estimation. By integrating a quantile-based cutting strategy, we improved the MDL computation time without compromising the performance. Experimental results on both synthetic and real datasets demonstrated the advantage of our approach compared to conventional uniform binning methods.
TABLE I: Models Performance in Multiclass classification of dry beans (mean ± std).